### v1

@@ Abstract
The so-called Geometric Trinity of Gravity includes General Relativity (GR), based on spacetime curvature; the Teleparallel Equivalent of GR (TEGR), which relies on spacetime torsion; and the Symmetric Teleparallel Equivalent of GR (STEGR), grounded in nonmetricity. Recent studies demonstrate that GR, TEGR, and STEGR are dynamically equivalent, raising questions about the fundamental structure of spacetime, the under-determination of these theories, and whether empirical distinctions among them are possible. The aim of this work is to show that they are equivalent in many features but not exactly in everything. In particular, their relationship with the Equivalence Principle (EP) is different. The EP is a deeply theory-laden assumption, which is assumed as fundamental in constructing GR, with significant implications for our understanding of spacetime. However, it introduces unresolved conceptual issues, including its impact on the nature of the metric and connection, its meaning at the quantum level, tensions with other fundamental interactions and new physics, and its role in dark matter and dark energy problems. In contrast, TEGR and STEGR recover the EP but do not rely on it as a foundational principle. The fact that GR, TEGR, and STEGR are equivalent in non-trivial predictions, but the EP is not necessary for TEGR and STEGR, suggests that it may not be a fundamental feature but an emergent one, potentially marking differences in the empirical content of the three theories. Thus, the developments within the Geometric Trinity framework challenge traditional assumptions about spacetime and may help to better understand some of the unresolved foundational difficulties related to the EP.
@@ 1 Introduction
Despite General Relativity (GR) being considered the ”standard theory” of gravity, there has never been a period in its history without serious alternatives being proposed and developed, with different approaches and motivations [1, 2, 3, 4]. Particularly in the last four decades, in the so-called ”precision cosmology era”, advancements in cosmological observations, gravitational wave physics and high precision tests have highlighted significant shortcomings, both at theoretical and observational level, both at small (UV) scales, and at large (IR) scales. [1].
@@
Several authors find it natural and elegant to consider improving the gravitational components of field equations [1, 2, 3], by exploring extended or alternative theories of gravity. There are many types of approaches to modified gravity theories [1, 2]: straightforward extensions like $f(R)$ gravity, Vector-Scalar-Tensor theories, where geometry can non-minimally couple to new fields; higher-order theories, where derivatives of metric components higher than second order can appear; theories with modified geometry; theories based on different principles, such as MOdified Newtonian Dynamics (MOND); unifying theories attempting to quantize gravity, and so on.
@@
A general approach can be inserted in the context of the so-called metric-affine theories of gravity, which fall into the category of approaches where geometry is enlarged and improved [5]. In particular, in 1919 Palatini showed that the metric tensor and the affine connection, which constitute GR, can be considered as two different geometric structures and can be varied independently [6]. These considerations led to the development of theories where the field equations can be formulated in terms of other geometric invariants: the torsion tensor and the non-metricity tensor. Together with the curvature tensor, considering these three geometric objects, we can build up different theories, where GR is a particular case in a lake of more general metric-affine theories, where torsion and non-metricity are set to zero. Among these theories, it is of particular interest the so-called Geometric Trinity of Gravity [7, 8, 5, 9], which comprises GR, built upon the metric tensor and grounded on the curvature of spacetime; the Teleparallel Equivalent of GR (TEGR), formulated in terms of torsion of spacetime and relying on tetrads and spin connection; and the Symmetric Teleparallel Equivalent of GR (STEGR), built on nonmetricity and constructed from metric tensor and affine connection. For this reason, in this case we can speak of modified spacetime rather than modified gravity, as some authors suggested [10].
@@
Significantly, these three theories have been found to be dynamically equivalent to GR, as the names suggest, and the Geometric Trinity recently gained a lot of attention in both the theoretical [7, 8, 5, 9, 11] and the philosophical literature [12, 13, 14, 15, 16]. From these recent theoretical developments, many questions arise. Should we consider TEGR and STEGR as proper alternative theories to GR or merely different dynamical formulations? Is it possible to empirically discriminate among them? If they are dynamically equivalent, is gravitation given by curvature, torsion, or non-metricity of spacetime at some fundamental level?
@@
In this work, it will be shown that a closer inspection on the equivalent features reveals crucial differences among the theories, that is, their relation with respect to the Equivalence Principle (EP). The EP is one of the most important assumptions of GR, which has been famously called as the ”midwife” who allowed Einstein to develop the theory [17, 18]. As we will see in detail, there are many open conceptual difficulties, which are direct or indirect consequences of the EP imposition at the foundation of the theory. Some of them puzzled Einstein himself until the very end of his life, such as the coincidence between the geodesic and the causal structure, or the fact that the fundamental object of the theory is the Riemannian metric $g_{\mu\nu}$ instead of the connection $\Gamma^{\rho}_{\mu\nu}$. The fact that the EP is such a theory-laden principle has led some authors to describe it as a ”beast” [18], or a bunch of beasts, given all its different formulations.
@@
Significantly, the Geometric Trinity suggests that there are viable theories of gravity which have not the necessity to impose the EP. The fact that these three representations of gravity are dynamically equivalent, and that the EP can be recovered but not at the foundation of TEGR and STEGR, suggests that the EP could be not a fundamental principle, but an emergent feature related to some symmetry or gauge [11, 19, 20]. Therefore, it will be argued that, at some level, the EP might constitute a direct difference in the empirical content between the three representations and could allow us to discriminate among them.
@@
This result is relevant because, if it is the case, it could allow to relax this theory-laden assumption at fundamental level and address some of the open problems. This, in some sense, is in line with the intuition by Synge, who considered it only a midwife and not a fundamental feature of the world [17].
@@ 2 General Relativity: assumptions and shortcomings
There are many significant conceptual difficulties that emerge from the assumption and the application of the EP, which are known in the physical and foundational literature, and which are sufficient to give to the EP the reputation of being ”a beast” [18].
Without entering in historical details, GR is built upon the EP, as well as upon other fundamental assumptions. There are various forms of the EP and many possible definitions for each form (see [18] for a foundational review and [4, 21, 22] for more theoretical and experimental reviews).
In its weaker form (WEP), the EP can be stated as the fact that all bodies fall in a gravitational field with the same acceleration, regardless of their mass or internal structure [4]. Then Einstein extended its scope, leading to the formulation of what is now known as the Einstein Equivalence Principle (EEP), which can be summarized as follows [4]:
  1. [1.]
  2. 1.
    The WEP is valid.
  3. 2.
    The outcome of any local non-gravitational experiment is independent of the velocity of the freely falling apparatus.
  4. 3.
    The outcome of any local non-gravitational experiment is independent of the location and time at which it is conducted in the universe.
Then, the definition of the EEP as the local validity of Special Relativity (SR) is now often called as the geometric formulation of the Strong EP (SEP). It can be stated as follows [4, 18]:
in any smooth four-dimensional manifold, it is possible to consider a small spacetime region where spatial and temporal gravitational changes are negligible. Therefore, there always exists a local inertial frame (LIF) where gravitational effects can be nullified.
In other words, inertial effects are locally indistinguishable from gravitational effects. Or to use the very words of Einstein from his definition of the EP [23]:
Gravity and inertia are the same in their very essence (‘wesensgleich’).
This link between the EEP and the local validity of SR was first underlined by Pauli in 1921 [18]. Where LIFs are defined by the Riemann theorem for every point $p\leavevmode\nobreak\ \in\leavevmode\nobreak\ M$, a manifold, in a local chart $(U,\phi)$ of $p$ as [5]:
$g_{\mu\nu}(\phi(p))=\eta_{\mu\nu},\;\;\;\nabla_{\lambda}g_{\mu\nu}=0\,.$ (1)
Note that Einstein did not call it as the SEP, but simply as the EP [18]. So he basically saw what we now call the EEP and the geometric formulation of the SEP as strictly related concepts.
@@
There are also other different formulations of the SEP, as the extension of the EEP also to bodies with non-negligible self-gravitational interactions and gravitational experiments [4] (see [18] for a discussion on the different definitions).
@@
In the following, we will list eight foundational problems relevant for this work. They are:
@@
As a consequence of the imposition of the SEP, the Christoffel symbols $\Gamma^{\rho}_{\mu\nu}$ coincide with the Levi-Civita connection, as we can see from the Levi-Civita theorem, which starts from the SEP [24]:
$\nabla_{\lambda}g_{\mu\nu}=0\rightarrow\Gamma^{\rho}_{\mu\nu}=\begin{Bmatrix}\rho\,\\ \mu\nu\end{Bmatrix}.$ (2)
Therefore, assuming the SEP, the unique possible affine symmetric connection is the Levi-Civita one, which contains the derivatives of $g_{\mu\nu}$:
$\Gamma^{\rho}_{\mu\gamma}=\begin{Bmatrix}\rho\,\\ \mu\nu\end{Bmatrix}=\frac{1}{2}g^{\rho\gamma}(\partial_{\mu}g_{\lambda\nu}+\partial_{\nu}g_{\mu\lambda}-\partial_{\lambda}g_{\mu\nu})\,.$ (3)
Consequently, by construction, the Levi-Civita connection has no dynamics, but it is a by-product of the metric $g_{\mu\nu}$. Physically, it represents the apparent forces acting on the body due to the curved geometric background. This means that the metric $g_{\mu\nu}$ determines, at the same time, the causal structure (light cones with rods and clocks) and the geodesic structure (the free fall of test particles) [24, 21]. However, it is important to underline that a priori there is no relation between the connection $\Gamma^{\rho}_{\mu\nu}$ and the metric tensor $g_{\mu\nu}$, but it is a consequence of the imposition of SEP. In fact, this coincidence does not work anymore for extensions of GR as $f(R)$ [25]. This unjustified coincidence is a first conceptual problem, which has been widely discussed in the literature (see e.g. [26]).
@@
The second problematic direct consequence of this picture is the following. As mentioned, in GR the real fundamental dynamical object is the metric tensor $g_{\mu\nu}$. However, it is not the gravitational field, but a set of potentials. The proper gravitational field is represented by the connection, as Einstein himself insisted many times [27, 18, 28]. It can be directly seen from the geodesic equation:
@@
where $\Gamma^{\rho}_{\mu\gamma}\frac{dx^{\mu}}{d\tau}\frac{dx^{\mu}}{d\tau}$ represents the generalization of the Newtonian forces.
This problem puzzled Einstein until the very end of his life, when he remarked again that the fundamental element should be the connection, and only indirectly the Riemannian metric $g_{\mu\nu}$ [29, pp. XVIII-XIX] (see also [30, p.9]). Moreover, in experiments, what we really measure are the forces (or the accelerations), which are represented by the connection $\Gamma^{\rho}_{\mu\nu}$. And as we have seen, the connections are the first $derivatives$ of the metric and the second derivatives of the local inertial coordinates:
@@
Therefore, thanks to the SEP, spacetime is described by the double $\{M,g_{\mu\nu}\}$, i.e. the Riemannian manifold, where $M$ is the manifold and $g_{\mu\nu}$ the metric tensor, that is the fundamental object.
@@
As a general consideration, it is possible to state that ”new physics” naturally predicts the violation of the EP at some level. The simplest case to observe a violation of EP, in the weaker form, is by Scalar-Tensor Theories, where, in the Einstein Frame, the mediation of the ”fifth force” causes a difference in the free fall between different objects [1]. Then, in theories featuring Quintessence, where the cosmological constant is replaced by a slowly evolving scalar field, one expects that the coupling of this field with matter induces gravitational forces that depend on the composition of the body. This clearly violates the EP. Scalar fields violating the EP are also predicted by theories involving extra dimensions, such as String Theory (ST). As argued in Ref.[31], current precision levels of WEP tests (today $10^{-15}$, as we will see in Sec. 5) should not discourage further research, as ST could imply WEP violations even further. For this reason, in Ref. [31], it is suggests that the WEP experiments are the most sensitive tools that we have to test new physics. Similarly, some authors argue that quantum properties of gravity could have observable experimental consequences at low energies, such as the dependence of geodesic motion on the mass of test particles, as explored in [32, 33]. Other authors [34] derive a direct violation of WEP at finite temperature from Quantum Field Theory. Then, famously, there is MOND (MOdified Newtonian Dynamics), which from observational constraints, should induce violations of the SEP [35]. Data from open clusters show a smaller mass discrepancy than would be required if MOND obey the SEP. Finally, it can be shown that MOND is just a particular case of some field theory (e.g. $f(R)$ gravity [36, 37]) so SEP can be questioned as soon as one relaxes the hypothesis that GR is the ”only” viable theory of gravity.
@@
iv) Validity of the EP at quantum level
A further significant issue is that we do not know if the EP is valid at quantum level. At the moment, we are only assuming its validity. We are not even sure if this principle could be generalized by the quantum formalism. There are attempts in this direction, but there are conflicting opinions among physicists [38, 21]. Since at quantum level particles behave like wave packets, it is difficult to make sense of the concepts of free fall universality or the identity between the gravitational and the inertial mass. Anyway, given quantum mechanics, we have no a priori reasons to postulate the EP validity.
@@
There is also the problem of Dark Energy (DE) and Cold Dark Matter (CDM), on which there is a big debate in the community. As already mentioned, several authors find more elegant to consider altering the gravitational component of the field equations [1] in order to address the dark phenomenology as a geometric problem, instead of a fluid one [39]. The main practical reason is that, up today, there is no final indication that dark side components could be addressed by new fundamental particle sector [40, 41, 42, 43]. This debate between the dark fluid hypothesis and modified gravity approaches is impressively widespread in the community of philosophers of physics [44, 3, 45, 46], with particular focus on the metric postulate. Authors argued that since any viable metric theory of gravity finds dark matter in the sieve [47], it could be the case that a non-metric theory of gravity could help in better understanding this problem.
The interesting thing is that, similarly to GR, where we can extend it to $f(R)$ gravity, $f(T)$ and $f(Q)$ gravity are the extensions of TEGR and STEGR, respectively. Where $R$ is the Ricci curvature scalar, $T$ is the torsion scalar and $Q$ the non-metricity scalar, while $f(R)$, $f(T)$ and $f(Q)$ are more general functions of them. The dynamical equivalence in the Geometric Trinity holds only for theories linear in the scalar invariants and not for the extensions, for different reasons [48, 5, 9]. Firstly, the extensions give rise to dynamics with different degrees of freedom. In particular, in $f(R)$ gravity, we have field equations of fourth order, in metric representation, whereas $f(T)$ and $f(Q)$ still remains of second-order. In addition, in $f(T)$ and $f(Q)$, we cannot choose, in general, a gauge to simplify the calculations, as in the cases of TEGR and STEGR. The point is that similarly to the fact that $f(R)$ theories are being studied to resolve shortcomings of GR at different scales (see [1]), also the extensions of TEGR and STEGR show interesting features in this direction. Physicists are already exploring $f(T)$ and $f(Q)$ gravities to study not only DE but also large structures, bouncing cosmologies, quantum cosmology, relativistic MOND theories, cosmography, inflation and gravitational waves (see for instance [48, 49]).
Therefore, the exploration of these alternatives to GR may provide a promising path to a successor and more fundamental theory. We can say ”more fundamental” because TEGR and STEGR have not to postulate the EP, as we will see, so they could be regarded as generalizations of GR, as Einstein already guessed.
@@
The EP sets gravity apart from other fundamental interactions of Nature. In today theoretical physics, it is believed to be very important to formulate theories as $gauge$ theories, since it works so well with other fundamental interactions [50]. The interesting fact is that gravitation can be reformulated as a gauge theory properly without assuming the EP in any form [5].

### v2

@@ Abstract
The so-called Geometric Trinity of Gravity includes General Relativity (GR), based on spacetime curvature; the Teleparallel Equivalent of GR (TEGR), which relies on spacetime torsion; and the Symmetric Teleparallel Equivalent of GR (STEGR), grounded in nonmetricity. Recent studies demonstrate that GR, TEGR, and STEGR are dynamically equivalent, raising questions about the fundamental structure of spacetime, the under-determination of these theories, and whether empirical distinctions among them are possible. The aim of this work is to show that they are equivalent in many features but not exactly in everything. In particular, their relationship with the Equivalence Principle (EP) is different. The EP is a deeply theory-laden assumption, which is assumed as fundamental in constructing GR, with significant implications for our understanding of spacetime. However, it introduces unresolved conceptual issues, including its impact on the nature of the metric and connection, its meaning at the quantum level, tensions with other fundamental interactions and new physics, and its role in dark matter and dark energy problems. In contrast, TEGR and STEGR recover the EP, in particular in its strong formulation, but do not rely on it as a foundational principle. The fact that GR, TEGR, and STEGR are equivalent in non-trivial predictions, but the EP is not necessary for TEGR and STEGR, suggests that it may not be a fundamental feature but an emergent one, potentially marking differences in the empirical content of the three theories. Thus, the developments within the Geometric Trinity framework challenge traditional assumptions about spacetime and may help to better understand some of the unresolved foundational difficulties related to the EP.
@@
Foundations of Physics (2025), 55:69
https://doi.org/10.1007/s10701-025-00882-x
Received: 19 January 2025 / Accepted: 28 July 2025
@@ 1 Introduction
Despite General Relativity (GR) being considered the “standard theory” of gravity, there has never been a period in its history without serious alternatives being proposed and developed, with different approaches and motivations [1, 2, 3, 4]. Particularly in the last four decades, in the so-called “precision cosmology era”, advancements in cosmological observations, gravitational wave physics and high precision tests have highlighted significant shortcomings, both at theoretical and observational level, both at small (UV) scales, and at large (IR) scales. [1].
@@
Several authors find it natural and elegant to consider improving the gravitational components of field equations [1, 2, 3], by exploring extended or alternative theories of gravity. There are many types of approaches to modified gravity theories [1, 2]: straightforward extensions like $f(\mathcal{R})$ gravity, Vector-Scalar-Tensor theories, where geometry can non-minimally couple to new fields; higher-order theories, where derivatives of metric components higher than second order can appear; theories with modified geometry; theories based on different principles, such as MOdified Newtonian Dynamics (MOND); unifying theories attempting to quantize gravity, and so on.
@@
A general approach can be inserted in the context of the so-called metric-affine theories of gravity, which fall into the category of approaches where geometry is enlarged and improved [5]. In particular, in 1919 Palatini showed that the metric tensor and the affine connection, which constitute GR, can be considered as two different geometric structures and can be varied independently [6]. These considerations led to the development of theories where the field equations can be formulated in terms of other geometric invariants: the torsion tensor and the non-metricity tensor. Together with the curvature tensor, considering these three geometric objects, we can build up different theories, where GR is a particular case in a lake of more general metric-affine theories, where torsion and non-metricity are set to zero. Among these theories, of particular interest is the so-called Geometric Trinity of Gravity [7, 8, 5, 9], which comprises GR, built upon the metric tensor and grounded on the curvature of spacetime; the Teleparallel Equivalent of GR (TEGR), formulated in terms of torsion of spacetime and relying on tetrads and spin connection; and the Symmetric Teleparallel Equivalent of GR (STEGR), built on nonmetricity and constructed from metric tensor and affine connection. For this reason, in this case we can speak of modified spacetime rather than modified gravity, as some authors suggested [10].
@@
Significantly, these three theories have been found to be dynamically equivalent to GR, as the names suggest, since their actions differ by boundary terms, which are dynamically irrelevant because the spacetime is asymptotically flat at infinity, and thus these terms are vanishing [5]. The Geometric Trinity recently gained a lot of attention in both the theoretical [7, 8, 5, 9, 11] and the philosophical literature [12, 13, 14, 15, 16]. From these recent theoretical developments, many questions arise. Should we consider TEGR and STEGR as proper alternative theories to GR or merely different dynamical formulations? Is it possible to empirically discriminate among them? If they are dynamically equivalent, is gravitation given by curvature, torsion, or non-metricity of spacetime at some fundamental level? Some authors have argued that TEGR is empirically equivalent to GR, having the same dynamics - that is, the same action up to a boundary term [17, 18]. However, as shown by other authors, the empirical content of a theory is not exhausted by its dynamics, i.e. by the equations of motion [19]. Therefore, two theories with the same action may still exhibit differences in their empirical content, as we shall argue.
@@
In this work, it will be shown that a closer inspection on the equivalent features reveals crucial differences among the theories, that is, their relation with respect to the Equivalence Principle (EP). The EP was one of the most important assumptions that led to the discovery of GR, famously described as the “midwife” that helped Einstein to develop the theory [20, 21]. As we will see in detail, there are many open conceptual difficulties, which are direct or indirect consequences of the EP imposition at the foundation of the theory. Some of them puzzled Einstein himself until the very end of his life, such as the coincidence between the geodesic and the causal structure, or the fact that the fundamental object of the theory is the Riemannian metric $g_{\mu\nu}$ instead of the connection $\Gamma^{\rho}_{\mu\nu}$. The fact that the EP is such a theory-laden principle has led some authors to describe it as a “beast” [21], or a bunch of beasts, given all its different formulations.
@@
Significantly, the Geometric Trinity suggests that there are viable theories of gravity which need not impose the EP. The fact that these three representations of gravity are dynamically equivalent, and that the EP can be recovered but not at the foundation of TEGR and STEGR, suggests that the EP could be not a fundamental principle, but an emergent feature related to some symmetry or gauge [11, 22, 23]. Therefore, it will be argued that, at some level, the EP might constitute a direct difference in the empirical content between the three representations and could allow us to discriminate among them.
@@
This result is relevant because, if it is the case, it could allow one to relax this theory-laden assumption at fundamental level and address some of the open problems. This, in some sense, is in line with the intuition by Synge, who considered it only a midwife and not a fundamental feature of the world [20].
In the EP discussion, we follow the approach by Lehmkuhl [21], as well as the recent contribution by Read and Teh on the EP in teleparallel gravity [24]. We adopt their framework and extend the discussion to the full Geometric Trinity.
@@ 2 General Relativity: assumptions and shortcomings
There are many significant conceptual difficulties that emerge from the assumption and the application of the EP, which are known in the physical and foundational literature, and which are sufficient to give to the EP the reputation of being “a beast” [21].
Without entering in historical details, GR is built upon the EP, as well as upon other fundamental assumptions. In the following, we adopt the Lehmkuhl framework [21], and we follow, in particular, the definitions provided by Read and Teh in a recent discussion of the EP in TEGR [24].
In its weaker form (WEP), the EP states that the gravitational mass of a body is equal to its inertial mass (called WEP2 in Ref. [21, p. 5] and [24, p. 3484]). The empirical indistinguishability between gravitational and inertial effects was a crucial starting point for Einstein in formulating his version of the EP, which is now known as the Einstein Equivalence Principle (EEP) [21, p. 10]:
EEP: Gravity and inertia are the same in their very essence (”wesensgleich”).
{justify}
Mathematically, it means that both gravitational and inertial effects are represented by the components of the same compatible connection [24, p. 3486].
Finally, the link between the EEP and the local validity of Special Relativity (SR), first underlined by Pauli in 1921 [21, pp. 21-22], constitutes the basis for the strong formulation of EP (SEP). According to the definition by Pauli, for every infinitely small region of the world, there always exists a coordinate system such that gravitational effects can be neglected. Einstein did not call it as the SEP, but he considered the EEP and the local validity of SR as two strictly related concepts [21]. Read and Teh state the SEP as follows [24, p. 3486]:
$\mathbf{SEP_{1,EEP}:}$ At any point $p\leavevmode\nobreak\ \in\leavevmode\nobreak\ M$, a manifold, one can find an orthonormal frame in which gravito-inertial effects, as represented by connection coefficients, vanish.
Then, Read and Teh define a version of the SEP which do not assume the EEP, which will be useful in relation to TEGR:
$\mathbf{SEP_{1,\neg EEP}:}$ At any point $p\leavevmode\nobreak\ \in\leavevmode\nobreak\ M$, a manifold, one can find a frame in which inertial effects, as represented by connection coefficients, cancel gravitational effects, as represented by some tensor quantity.
Adding that the frames in which the principle holds are related by Lorentz transformations, we arrive to what Read and Teh call $SEP_{2}$ [24, p. 3487]:
$\mathbf{SEP_{2,EEP}:}$ SEP1,EEP holds, and the frames in which this principle holds are related by Lorentz transformations.
$\mathbf{SEP_{2,\neg EEP}:}$ SEP1,¬EEP holds, and the frames in which this principle holds are related by Lorentz transformations.
The standard frame where the physics of SR is locally recovered, the metric tensor $g_{\mu\nu}$ takes the Minkowski form $\eta_{\mu\nu}$, and the connection coefficients $\Gamma^{\rho}_{\mu\nu}$ vanish, is called the Local Inertial Frame (LIF). The requirement that such a frame exists at every point $p\leavevmode\nobreak\ \in\leavevmode\nobreak\ M$, demands the metric compatibility condition $\nabla_{\lambda}g_{\mu\nu}=0$ [25, p. 313].
@@
Note that there are also other different formulations of the SEP, as the extension of the EEP also to bodies with non-negligible self-gravitational interactions and gravitational experiments [4, p. 76]. However, the relation between these different definitions is not yet clear and this issue is not addressed in this work.
@@
In the following, we will list eight foundational problems relevant for this work.
@@
As a consequence of the imposition of the SEP, the Christoffel symbols $\Gamma^{\rho}_{\mu\nu}$ coincide with the Levi-Civita connection. Assuming the SEP, the unique possible affine symmetric and metric compatible connection is the Levi-Civita one. The a priori assumption of the SEP in this way selects the Levi-Civita connection [25, p. 314]:
$g_{\mu\nu}(p)=\eta_{\mu\nu},\quad\Gamma^{\rho}_{\mu\nu}(p)=0,\quad\nabla_{\lambda}g_{\mu\nu}=0,\quad\Gamma^{\rho}_{\mu\nu}=\Gamma^{\rho}_{\nu\mu}$ (1)
$\Downarrow$
$\Gamma^{\rho}_{\mu\gamma}=\begin{Bmatrix}\rho\,\\ \mu\nu\end{Bmatrix}=\frac{1}{2}g^{\rho\gamma}\left(\partial_{\mu}g_{\lambda\nu}+\partial_{\nu}g_{\mu\lambda}-\partial_{\lambda}g_{\mu\nu}\right)$ (2)
Consequently, by construction, the Levi-Civita connection has no indipendent dynamics, but it is a by-product of the metric $g_{\mu\nu}$, containing its derivatives. Physically, it represents the apparent forces acting on the body due to the curved geometric background. This means that the metric $g_{\mu\nu}$ determines, at the same time, the causal structure (light cones with rods and clocks) and the geodesic structure (the free fall of test particles) [25, 26]. However, it is important to underline that a priori there is no relation between the connection $\Gamma^{\rho}_{\mu\nu}$ and the metric tensor $g_{\mu\nu}$, but it is a consequence of the imposition of SEP. In fact, this coincidence does not work anymore for extensions of GR as $f(\mathcal{R})$ [27]. This unjustified coincidence is a first conceptual problem, which has been widely discussed in the literature (see e.g. [28]).
@@
The second problematic direct consequence of this picture is the following. As mentioned, in GR the truly fundamental dynamical object is the metric tensor $g_{\mu\nu}$. However, there are reasons not to regard it as the gravitational field, since it represents a set of potentials. According to Einstein himself, the proper gravitational field should be identified with the connection components [29, 30, 31, 32], as they are derivatives of the metric and represent a straightforward generalization of the Newtonian gravitational field, which is the gradient of the potential [31, p. 846]. The connection components also appear in the geodesic equation, which Einstein considered the successor and generalization of the Newtonian law of inertia [21, p. 15], thus making them natural candidates for representing the gravitational field:
@@
where $\Gamma^{\rho}_{\mu\gamma}\frac{dx^{\mu}}{d\tau}\frac{dx^{\mu}}{d\tau}$ represents the generalization of the Newtonian forces. For a discussion of other candidates for the gravitational field, see Ref. [30].
This problem puzzled Einstein until the very end of his life, when he remarked again that the fundamental element should be the connection, and only indirectly the Riemannian metric $g_{\mu\nu}$ [33, pp. XVIII-XIX] (see also [34, p. 9]). Moreover, in experiments, what we really measure are the forces (or the accelerations), which are represented by the connection $\Gamma^{\rho}_{\mu\nu}$. And as we have seen, the connections are the first derivativesderivatives of the metric and the second derivatives of the local inertial coordinates:
@@
Therefore, thanks to the SEP, spacetime is described by the double $\langle M,g_{\mu\nu}\rangle$, i.e. the Riemannian manifold, where $M$ is the manifold and $g_{\mu\nu}$ the metric tensor, that is the fundamental object.
@@
As a general consideration, it is possible to state that “new physics” naturally predicts the violation of the SEP at some level. The simplest case is in Scalar-Tensor Theories, where, in the Einstein Frame, the mediation of the “fifth force” causes a difference in the free fall between different objects [1], which violates the WEP, and therefore the SEP. Then, in theories featuring Quintessence, where the cosmological constant is replaced by a slowly evolving scalar field, one expects that the coupling of this field with matter induces gravitational forces that depend on the composition of the body. This clearly violates the WEP. Scalar fields violating the WEP are also predicted by theories involving extra dimensions, such as String Theory (ST). As argued in Ref.[35], current precision levels of WEP tests (today $10^{-15}$, as we will see in Sec. 5) should not discourage further research, as ST could imply WEP violations even further. For this reason, in Ref. [35], it is suggests that the WEP experiments are the most sensitive tools that we have to test new physics. Similarly, some authors argue that quantum properties of gravity could have observable experimental consequences at low energies, such as the dependence of geodesic motion on the mass of test particles, as explored in [36, 37]. Other authors [38] derive a direct violation of WEP at finite temperature from Quantum Field Theory. Then, famously, there is MOND (MOdified Newtonian Dynamics), which from observational constraints, should induce violations of the SEP [39]. Data from open clusters show a smaller mass discrepancy than would be required if MOND obey the SEP, even though it seems that these predictions consider the definition of SEP as the extension of EEP also to bodies with non-negligible self-gravitational interactions [4, p. 76], which is not the same concept under consideration in this work. Finally, it can be shown that MOND is just a particular case of some field theory (e.g. $f(\mathcal{R})$ gravity [40, 41]) so SEP can be questioned as soon as one relaxes the hypothesis that GR is the “only” viable theory of gravity.
@@
iv) Validity of the SEP at quantum level
A further significant issue is that we do not know if the WEP, and therefore the SEP is valid at quantum level. At the moment, we are only assuming its validity. We are not even sure if this principle could be generalized by the quantum formalism. There are attempts in this direction, but there are conflicting opinions among physicists [42, 26, 43]. Since at quantum level particles behave like wave packets, it is difficult to make sense of the concepts of free fall universality or the identity between the gravitational and the inertial mass. Anyway, given quantum mechanics, we have no a priori reasons to postulate the WEP validity.
@@
There is also the problem of Dark Energy (DE) and Cold Dark Matter (CDM), on which there is a big debate in the community. As already mentioned, several authors find it more elegant to consider altering the gravitational component of the field equations [1] in order to address the dark phenomenology as a geometric problem, instead of a fluid one [44]. The main practical reason is that, up to today, there is no final indication that dark side components could be addressed by new fundamental particle sector [45, 46, 47, 48]. This debate between the dark fluid hypothesis and modified gravity approaches is impressively widespread in the community of philosophers of physics [49, 3, 50, 51], with particular focus on the metric postulate. Authors argued that since any viable metric theory of gravity finds dark matter in the sieve [52], it could be the case that a non-metric theory of gravity could help in better understanding this problem.
The interesting thing is that, similarly to GR, where we can extend it to $f(\mathcal{R})$ gravity, $f(T)$ and $f(Q)$ gravity are the extensions of TEGR and STEGR, respectively. Where $R$ is the Ricci curvature scalar, $T$ is the torsion scalar and $Q$ the non-metricity scalar, while $f(\mathcal{R})$, $f(T)$ and $f(Q)$ are more general functions of them. The dynamical equivalence in the Geometric Trinity holds only for theories linear in the scalar invariants and not for the extensions, for different reasons [53, 5, 9]. Firstly, the extensions give rise to dynamics with different degrees of freedom. In particular, in $f(\mathcal{R})$ gravity, we have field equations of fourth order, in metric representation, whereas $f(T)$ and $f(Q)$ still remains of second-order. In addition, in $f(T)$ and $f(Q)$, we cannot choose, in general, a gauge to simplify the calculations, as in the cases of TEGR and STEGR. The point is that similarly to the fact that $f(\mathcal{R})$ theories are being studied to resolve shortcomings of GR at different scales (see [1]), also the extensions of TEGR and STEGR show interesting features in this direction. Physicists are already exploring $f(T)$ and $f(Q)$ gravities to study not only DE but also large structures, bouncing cosmologies, quantum cosmology, relativistic MOND theories, cosmography, inflation and gravitational waves (see for instance [53, 54]).
Therefore, the exploration of these alternatives to GR may provide a promising path to a successor and more fundamental theory. We can say “more fundamental” because TEGR and STEGR have not to postulate the EP in any form, as we will see, so they could be regarded as generalizations of GR, as Einstein already guessed.
@@
The SEP sets gravity apart from other fundamental interactions of Nature. In today’s theoretical physics, it is believed to be very important to formulate theories as gaugegauge theories, since it works so well with other fundamental interactions [55]. The interesting fact is that gravitation can be reformulated as a gauge theory properly without assuming the EP in any form [5]. See Ref. [18, 56, 57, 58, 59] for philosophical discussions on TEGR and STEGR as gauge theories.